OpenStudy (blaycake):

Will reward with a medal! Set up the following word problem as a proportion and solve. Henry drove 65 miles on 3 gallons of gas. How far he can travel on a full 16 gallon tank? Set up a proportion.

OpenStudy (blaycake):

I just need it set up as a number problem, so i can solve the rest myself :) please and thank you

OpenStudy (kainui):

Make a guess, I think that'll be a good starting point.

OpenStudy (blaycake):

65+3n=x? I really am not sure how to set up the problem :/

OpenStudy (blaycake):

Hmm, 16 gallons so it would have to equal 16 right so i have to be wrong on that. and you can get 65 miles on 3 gallons

OpenStudy (blaycake):

I have to have this extra credit turned in about a half hour and this is the last question, can someone please help me set this up so i can finish the series of questions that goes with this once i have it set up

sammixboo (sammixboo):

Try to figure out how far you can go with 1 gallon

OpenStudy (blaycake):

20.6

OpenStudy (blaycake):

so it would be 20.6 times 16? im not trying to solve it at the moment, i have no problem solving it, i need to set it up like the word problem

OpenStudy (blaycake):

im just not sure how to set it up like that

OpenStudy (kainui):

So the way you can think about division is "something per something" so here they say: "Henry drove 65 miles on 3 gallons of gas." Which is the same as saying Henry went 65 miles per 3 gallons of gas, which is written as \(\frac{65}{3}\). "How far he can travel on a full 16 gallon tank?" This proportion will be the same since the more gas you have the more miles you go, so we'll end up with "Henry drives x miles per 16 gallons" which we can write as \(\frac{x}{16}\) and these two fractions are equal, and you can solve for x, the distance he can go.

sammixboo (sammixboo):

\(\tt \dfrac{65}{3}=\dfrac{x}{16}\)

OpenStudy (mathmale):

The whole point of this problem is to use ratios and equations of ratios correctly. Please make a point of looking for examples online or in your study material of how to do this. @sammixboo's equation is an example of the way to go. Note that the number of gallons, in each case, is in the denominator, and the number of miles in the numerator. Obviously, if this driver can go 65 miles on 3 gallons of fuel, he or she can go a lot further on 16 gallons. Use cross multiplication to solve this equation for x. x represents the distance the driver can go on 16 gallons.